OpenStudy (anonymous):

5. 8 = 3(3x + 8) − x how do i work this when combing like terms

OpenStudy (texaschic101):

do you know how to distribute through the parenthesis ?

OpenStudy (texaschic101):

basically, you take the number right before the parenthesis, and multiply it by everything in the parenthesis...this will get rid of the parenthesis. 5.8 = 3 * 3x + 3 * 8 5.8 = 9x + 24 --- now subtract 24 from both sides 5.8 - 24 = 9x -18.2 = 9x -- now divide both sides by 9 -18.2/9 = x -2.02 = x wait...is the problem 5.8 = ....or is it just 8 = and the 5 is the number of the problem ?

OpenStudy (texaschic101):

is this the whole problem...because I see a lot of black diamonds in your problem

OpenStudy (anonymous):

its 5.8=3(3x+8)-x @texaschic101

OpenStudy (texaschic101):

sorry...let me do this again 5.8 = 3(3x + 8) - x --- distribute through the parenthesis 5.8 = 9x + 24 - x --- subtract 24 from both sides 5.8 - 24 = 9x - x --- combine like terms -18.2 = 8x --- divide both sides by 8 -18.2/8 = x -2.275 = x check... 5.8 = 3(3(-2.275) + 8) - (-2.275) 5.8 = 3(-6.825 + 8) + 2.275 5.8 = 3(1.175) + 2.275 5.8 = 3.525 + 2.275 5.8 = 5.8 (correct) so x = - 2.275

OpenStudy (anonymous):

thank you @texaschic101